Staff:

LBTO: Juan Carlos

Observer: Shark_nir team (remote)

Summary.

Details.

09:00 am

contacted to help with Shark. Not one(AO Support) was schedule to support this activity

Argos cal unit not responding.
  • controller in "funny state"
  • reset controller from controller interface.
  • load the working parameters.
Argos cal unit in operations.

hexapod positions for PSF.

0.8,0.6,1.66,136,-591,0

09:52

Loop closed.

gain:LNgain

AOS status -> StartAO.

Shark-Nir pupil aligned.

Send the shark script to align the pupil, script didn;t do anything in WFs side (not offsets) and the script finished wiyhout an error.

Shark-nir team debuging.

The script computed some offsets to apply, moving the stages manually with the delta offsets.

x=0.835, y=-2.231

taking shark image to check if the offset was correct. And they are correct....

10:55

Shark-Nir still debugging.

Loop closed still closed.

12:14PM

still close loop.

Shark nir doing their stuff.

asked to the telescope how much time Shark_nir can be at the telescope; Answer until 4:30PM

passed info to shark-nir team.

2:30PM

StopAO

argos parked.light off.

WFs poarked.

Shark parker arm.

-- JuanCarlosGuerra - 15 Sep 2023
Topic revision: r3 - 15 Sep 2023, JuanCarlosGuerra
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback